You are financing a new car with a 3 year loan at 5% annual interest, compounded
monthly. The amount you are borrowing is $16,000. What are your monthly payments?
Answer #1

Borrowed amount (P) = $16000

Annual Interest = 5% = 0.05

Since, it is monthly compounded, monthly interest = Annual interest/12

Monthly Interest (R) = 0.05/ 12 = 0.00417

No.of payments = 3 years * 12 months

Therefore, no.of monthly payments (N) = 36

Calculation of EMI:

EMI = P * R * \frac{(1+R)^{N}}{(1+R)^{N}-1}

EMI = 16000 * 0.00417 * \frac{(1+0.00417)^{36}}{(1+0.00417)^{36}-1}

EMI = 66.72 * \frac{(1.162)}{(1.162)-1}

EMI = 66.72 * \frac{1.162}{0.162}

EMI = 66.72 * 7.188

EMI = 479.563

Therefore, the monthly payment is $479.563